An average human weights about 650 N.if two such generic humans each carried 1.0 coulomb of excess charge,one positive and one negative. How far apart would they have to be for the electric attraction between them to equal their 650-N weight


1
Expert's answer
2022-03-10T10:17:35-0500

Answer

So the distance is given by

"R=\sqrt{\frac{kqq'}{F}}\\=\sqrt{\frac{9*1*1}{650}}\\=11.74cm"